OpenStudy (anonymous):

The equation of a curve is given by y= 1/6 (2x - 9)^6 . a) Find the x coordinate of the point at which the tangent to the curve is parallel to the x- axis. b) Find the equation of tangent to the curve at x = 5 c) The normal to the curve at x = 5 meets the y axis at point A. Find the coordinates of point A. d) A point S (x,y) moves along the curve with time t. When y = 1/6 , dy/dt = 4 units/s, find the corresponding rate of change of x.

OpenStudy (anonymous):

Find the x coordinate of the point at which the tangent to the curve is parallel to the x- axis. take the derivative, set it equal to zero and solve for \(x\)

OpenStudy (anonymous):

you got that? you can almost do it in your head

OpenStudy (anonymous):

I got x = 4.5

OpenStudy (anonymous):

\[y'=(2x-9)^5\] so it is zero at \(x=\frac{9}{2}=4.5\) yes

OpenStudy (anonymous):

Alright thnx

OpenStudy (anonymous):

Find the equation of tangent to the curve at x = 5 put \(x=5\) in the derivative to find the slope

OpenStudy (anonymous):

\[m=(2\times 5-9)^5=1^5=1\]

OpenStudy (anonymous):

then use the point - slope formula with \(m=1, x_1=5, y_1=\frac{1}{6}\)

OpenStudy (anonymous):

Okay i got y = 2x - 9/5 is that right

OpenStudy (anonymous):

can't be because the slope is 1

OpenStudy (anonymous):

\[y-\frac{1}{6}=x-5\]

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Hw about part c and d

OpenStudy (anonymous):

The derivative is 2(2x-9)^5 i am sure about it